# Open-Set 3D Semantic Instance Maps for Vision Language Navigation - O3D-SIM

### v1

@@ Abstract
vision-language-navigation, semantic, 3D-representation, chatgpt, RGB-D perception
@@ 1 Introduction
Building upon [1], we introduce a novel approach for Semantic Instance Maps. Our work introduces Open-set 3D Semantic Instance Maps (O3D-SIM), addressing the limitation of traditional closed-set methods [7, 8], which assume only predefined set objects will be encountered during operation. In contrast, O3D-SIM leverages methods to identify and potentially categorize unseen objects not explicitly included in the training data. This is crucial for real-world scenarios with diverse and unseen objects. This enables better performance and complex query handling for unseen object categories. We address the previously mentioned limitations by enabling instance-level object identification within the spatial representation and operating in an open-set manner, excelling in real-world scenarios. We have achieved major improvements over our previous work, which operated in a closed-set manner. The current pipeline leverages state-of-the-art foundational models like CLIP [9] and DINO [10]. These models extract semantic features from images, allowing them to recognize objects and understand the finer details and relationships between them. For example, the DINOv2 [10] model, trained on various chair images, can identify a chair in a new image and distinguish between a dining chair and an office chair.
@@
The remainder of this paper is organized as follows. Section 2 reviews and analyzes recent literature, providing the necessary background on semantic scene understanding, 3D scene reconstruction, and VLN to the reader. Section 3 outlines the methodology behind our proposed 3D map. Section 4 discusses the experimental evaluation of our proposed 3D map’s effectiveness. Finally, concluding remarks and directions for future work are presented in Section 5.
@@ 2.1 Vision-and-Language Navigation:
A major challenge in VLN is interpreting language instructions in unfamiliar environments. A significant limitation of previous studies in this domain is their handling of action errors. If a robot agent makes an incorrect action, it risks failing to reach its destination or exploring unnecessary areas, leading to increased computational demands and possibly entering a state from which recovery is unfeasible. State-of-the-art VLN methodologies employ diverse strategies to excel in such scenarios. Some methods adopt a specialized pre-training and fine-tuning approach designed explicitly for VLN tasks, utilizing transformer-based architectures [13]. These strategies often involve using image-text-action triplets in a self-supervised learning context [14]. Other approaches refine the pre-training process to enhance VLN task performance, for instance, by emphasizing the learning of spatiotemporal visual-textual relationships to better utilize past observations for future action prediction [15, 16, 17]. Furthermore, contemporary VLN systems predominantly rely on simulations due to their dependency on panoramic views and extracting region features, which can be computationally prohibitive. In contrast, our work demonstrates our pipeline’s efficiency and computational viability with real-world data, underscoring its practical applicability.
@@ 3 Methodology
The pipeline of the O3D-SIM creation is depicted in Fig.2. The first step of the creation of the O3D-SIM, presented in Section 3.2, is the extraction of the open-set semantic instance information from the RGB sequence of input images. This information includes, for each object instance, the mask information and the semantic features represented by the CLIP [9] and DINO [10] embedding features. The second step, presented in Section 3.3, uses this open-set semantic instance information to cluster the input 3D point cloud into an open-set semantic 3D objects map, see Figures 2 and 3. The operation is improved incrementally by applying the sequence of RGB-D images over time.
@@ 3.2 Open-set Semantic Information from Images
Figure 2: An overview of the proposed 3D mapping pipeline. Labels generated by the RAM model are input into Grounding DINO to generate bounding boxes for the detected labels. Subsequently, instance masks are created using the SAM model, while CLIP and DINOv2 embeddings are extracted in parallel. These masks, along with the semantic embeddings, are back-projected into 3D space to identify 3D instances. These instances are then refined using a density-based clustering algorithm to produce the O3D-SIM.
@@ 3.2.1 Open-set Semantic and Instance Masks Detection
The RAM model) [33] processes the input RGB image to produce semantic labelling of the object detected in the image. It is a robust foundational model for image tagging, showcasing remarkable zero-shot capability in accurately identifying various common categories. The output of this model associates every input image with a set of labels that describe the object categories in the image. The process begins with accessing the input image and converting it to the RGB colour space, then resized to fit the model’s input requirements, and finally transforming it into a tensor, making it compatible with the analysis by the model. Following this, the RAM model generates labels, or tags, that describe the various objects or features present within the image. A filtration process is employed to refine the generated labels, which involves the removal of unwanted classes from these labels. Specifically, irrelevant tags such as ”wall”, ”floor”, ”ceiling”, and ”office” are discarded, along with other predefined classes deemed unnecessary for the context of the study. Additionally, this stage allows for the augmentation of the label set with any required classes not initially detected by the RAM model. Finally, all pertinent information is aggregated into a structured format. Specifically, each image is catalogued within the img__\__dict dictionary, which records the image’s path alongside the set of generated labels, thus ensuring an accessible repository of data for subsequent analysis.
@@ 3.3 Creating the Open-set 3D Representation
Figure 3: Clustering in 3D of object instances using semantic and geometric features. Semantic similarity is verified using CLIP and DINOv2 embeddings. Geometric similarity is verified using 3D bounding boxes and overlap matrices.
@@ 3.3.1 The O3D-SIM Intialization
The set of object’s point clouds is processed further using DBSCAN[34] clustering for representation refinement. The point cloud is downsampled via voxel grid filtering to reduce the number of points and the computational complexity while preserving the data spatial structure manageable. DBSCAN groups points that are closely packed together while labelling points that lie alone in low-density regions as noise. In a post-clustering step, the largest cluster typically corresponds to the main object of interest within the point cloud is identified. This helps filter out the noise and irrelevant points, producing a cleaner representation of the object of interest.
@@ 3.3.2 Incremental Update of the O3D-SIM
Objects are detected for each new image, and new nodes are created like the initial image. These temporary nodes contain the 3D data for newly detected objects that must either be merged into the existing scene or added as new nodes. The similarity between newly detected and existing scene nodes is determined by combining visual similarity, derived from feature embeddings, and spatial (geometric) similarity, obtained from point cloud overlap, to formulate an aggregate similarity measure. If this measure surpasses a predetermined threshold, the new detection is deemed to correspond to an existing object in the scene. Indeed, the newly detected node is either merged with an existing scene node or added as a new node.
@@ 3.4 Language-Guided Navigation
In this section, we leverage the LLM-based approach from [1], which uses ChatGPT [35] to understand and map language commands to pre-defined function primitives that the robot can understand and execute. However, there are a few differences between our current approach and the approach in [1] regarding the use case of the LLM and the implementation of our function primitives. The previous approach used the LLM’s ability to bring in an open-set understanding by mapping general queries to the already-known closed-set class labels obtained via Mask2Former [7]. However, given the open-set nature of our new representation, O3D-SIM, the LLM does not need to do that. Figure  4 shows both approaches’ code output differences. The function primitives work similarly to the older approach, requiring the desired object type and its instance as an input. But now, the desired object is not from a pre-defined set of classes but a small query defining the object, so the implementation to find the desired location changes. We use the text and image-aligned nature of CLIP embeddings to find the desired object, where the input description is passed to the model, and its corresponding embedding is used to find the object in O3D-SIM. A cosine similarity is calculated between the embedding of the description and all the embeddings of our representation. These are ranked in a decreasing order, and the desired instance is selected. Once the instance is finalized, a goal corresponding to this instance is generated and passed to the navigation stack for autonomous navigation of the robot, hence achieving Language-Guided Navigation.
@@ 4 Experiments
Having introduced the O3D-SIM creation pipeline and its integration with ChatGPT for natural language understanding and Vision-Language Navigation (VLN) enhancement, we now turn to the evaluation of this novel representation both quantitatively and qualitatively. This will also shed light on the impact of the O3D-SIM representation on an agent’s ability to execute queries that mimic human interaction. The evaluation is structured into two subsections: Section 4.1 focuses on the quantitative assessment of O3D-SIM, and Section 4.2 addresses the qualitative analysis of the representation.
@@ 4.1 Quantitative Evaluation
To facilitate the construction of O3D-SIM and its quantitative evaluation, we employ the Matterport3D dataset [36] within the Habitat simulator [37]. Matterport3D, a comprehensive RGB-D dataset, encompasses 10,800 panoramic views derived from 194,400 RGB-D images across 90 large-scale buildings. It offers surface reconstructions, camera poses, and 2D and 3D semantic segmentations — critical components for creating accurate Ground-truth models. Both Matterport3D and Habitat are widely utilized for assessing the navigational abilities of VLN agents in indoor settings, enabling robots to execute navigational tasks dictated by natural language commands in a seamless environment, with performance meticulously documented. To evaluate O3D-SIM, we compiled 5,267 RGB-D frames and their respective pose data from five distinct scenes, applying this dataset across all mapping pipelines included in our assessment. Additionally, we gathered real-world environment data for evaluation purposes, thereby expanding our analysis to encompass six unique scenes.
@@
Results: We present the results of the evaluation metric Success Rate in Table 1. In our experimentation, we observe a remarkable improvement in performance compared to the other approaches we have shown in our paper, especially against the baselines from [1]. O3D-SIM performs better than VLMaps with CC due to its ability to identify instances robustly. SI Maps and O3D-SIM perform better than the baselines due to their ability to separate instances. However, O3D-SIM has the edge over SI Maps due to its open set and 3D nature, allowing it to understand the surroundings better.
@@
Table 1: O3D-SIM outperforms the baseline methods from [1] by significantly large margins on the Success Rate metric. It also shows an improvement from our previous approach, i.e., SI-Maps. The best results are highlighted in bold. For this evaluation, the agent executes a set of open-set and closed-set queries in 5 different scenes from Matterport3D and 1 real word environment.
@@ 4.2 Qualitative Results
Our approach excels in recognizing instance-level semantics, accurately identifying 5 out of 6 table instances (with one false positive), underscoring its precision across both simulated and real-world data. This demonstrates the robustness of our pipeline, further evidenced by the clarity of the semantic map and the ease with which instance-level segmentation results can be visualized. While methods like VLMaps might identify a broader range of objects due to their open-set nature, and SI-Maps may detect multiple instances of the same object, O3D-SIM uniquely excels at both, offering a comprehensive solution.
Figure 5: This figure presents the qualitative results for O3D-SIM. Mapping sequence 1 was recorded using a Kinect Azure on a mobile robot, whereas mapping sequence 2 was recorded using an iPad, both in the real world. The environment was changed between the two sequences.

### v2

@@ Abstract
Project Page - https://smart-wheelchair-rrc.github.io/o3d-sim-webpage
@@
vision-language-navigation, semantic-instance-segmentation, 3D-representation, ChatGPT, LLMs, RGB-D perception
This article is published in Taylor & Francis Advanced Robotics. DOI - 10.1080/01691864.2024.2395926
@@ 1 Introduction
Building upon [1], we introduce a novel approach for Semantic Instance Maps. Our work introduces Open-set 3D Semantic Instance Maps (O3D-SIM), addressing the limitation of traditional closed-set methods [7, 8], which assume only predefined set objects will be encountered during operation. In contrast, O3D-SIM leverages methods to identify and potentially categorize unseen objects not explicitly included in the training data. This is crucial for scenarios with diverse and unseen objects. This enables better performance and complex query handling for unseen object categories. We address the previously mentioned limitations by enabling instance-level object identification within the spatial representation and operating in an open-set manner, excelling in real-world scenarios. We have achieved major improvements over our previous work, which operated in a closed-set manner. The current pipeline leverages state-of-the-art foundational models like CLIP [9] and DINO [10]. These models extract semantic features from images, allowing them to recognize objects and understand the finer details and relationships between them. For example, the DINOv2 [10] model, trained on various chair images, can identify a chair in a new image and distinguish between a dining chair and an office chair.
@@
The remainder of this paper is organized as follows. Section 2 reviews and analyzes recent literature, providing the necessary background on semantic scene understanding, 3D scene reconstruction, and VLN to the reader. Section 3 outlines the methodology behind our proposed 3D map. Section 4 discusses the experimental evaluation of our proposed 3D map’s effectiveness in achieving complex language queries as shown in Figure 1. Finally, concluding remarks and directions for future work are presented in Section 5.
@@ 2.1 Vision-and-Language Navigation:
A major challenge in VLN is interpreting language instructions in unfamiliar environments. A significant limitation of previous studies in this domain is their handling of action errors. Suppose a robot agent makes an incorrect action. In that case, it risks failing to reach its destination or exploring unnecessary areas, leading to increased computational demands and possibly entering a state from which recovery is unfeasible. State-of-the-art VLN methodologies employ diverse strategies to excel in such scenarios. Some methods adopt a specialized pre-training and fine-tuning approach designed explicitly for VLN tasks, utilizing transformer-based architectures [13]. These strategies often involve using image-text-action triplets in a self-supervised learning context [14]. Other approaches refine the pre-training process to enhance VLN task performance, for instance, by emphasizing the learning of spatiotemporal visual-textual relationships to better utilize past observations for future action prediction [15, 16, 17].
In this work, we also address some of the limitations observed in existing end-to-end VLN methods, such as those detailed in Qi et al. [18], Huo et al. [19], and Hong et al. [20]. These approaches rely on direct action estimation from multimodal inputs without constructing an environmental map, necessitating highly specific navigational instructions, which can feel unnatural and rigid in real-world settings. Furthermore, these methods modify BERT [21] architectures to parse language commands, thus missing out on the broader contextual understanding offered by newer larger language models (LLMs) like GPT 3.5. These models also tend to process multiple RGB images per timestep to determine actions, leading to higher computational demands during inference. In contrast, our method integrates environmental mapping, which enhances the system’s spatial understanding and allows for more context-aware navigation.
Furthermore, contemporary VLN systems predominantly rely on simulations due to their dependency on panoramic views and extracting region features, which can be computationally prohibitive. In contrast, our work demonstrates our pipeline’s efficiency and computational viability with real-world data, underscoring its practical applicability.
@@
Other approaches to VLN, such as [1, 2, 4, 22], involve building a map of the environment along with the semantic understanding and utilising this global-level representation to achieve multimodal tasks. These works have shown that understanding the environment on a global scale helps achieve VLN tasks easily and saves the agent from unnecessary explorations that may occur due to wrong future-action predictions. Our approach falls under the same category.
@@ 3 Methodology
The pipeline of the O3D-SIM creation is depicted in Figure 2. The first step of the creation of the O3D-SIM, presented in Section 3.2, is the extraction of the open-set semantic instance information from the RGB sequence of input images. This information includes, for each object instance, the mask information and the semantic features represented by the CLIP [9] and DINO [10] embedding features. The second step, presented in Section 3.3, uses this open-set semantic instance information to cluster the input 3D point cloud into an open-set semantic 3D objects map, see Figures 2 and 3. The operation is improved incrementally by applying the sequence of RGB-D images over time.
@@ 3.2 Open-set Semantic Information from Images
Figure 2: An overview of the proposed 3D mapping pipeline. Labels generated by the RAM model are input into Grounding DINO to generate bounding boxes for the detected labels. Subsequently, instance masks are created using the SAM model, while CLIP and DINOv2 embeddings are extracted in parallel. These masks, along with the semantic embeddings, are back-projected into 3D space to identify 3D instances. These instances are then refined using a density-based clustering algorithm to produce the O3D-SIM. Figure3 shows how the object instance clustering works as part of O3D-SIM.
@@ 3.2.1 Open-set Semantic and Instance Masks Detection
The RAM model) [37] processes the input RGB image to produce semantic labelling of the object detected in the image. It is a robust foundational model for image tagging, showcasing remarkable zero-shot capability in accurately identifying various common categories. The output of this model associates every input image with a set of labels that describe the object categories in the image. The process begins with accessing the input image and converting it to the RGB colour space, then resized to fit the model’s input requirements, and finally transforming it into a tensor, making it compatible with the analysis by the model. Following this, the RAM model generates labels, or tags, that describe the various objects or features present within the image. A filtration process is employed to refine the generated labels, which involves the removal of unwanted classes from these labels. Specifically, irrelevant tags such as ”wall”, ”floor”, ”ceiling”, and ”office” are discarded, along with other predefined classes deemed unnecessary for the context of the study. Additionally, this stage allows for the augmentation of the label set with any required classes not initially detected by the RAM model. Finally, all pertinent information is aggregated into a structured format. Specifically, each image is catalogued within the img_\_dict dictionary, which records the image’s path alongside the set of generated labels, thus ensuring an accessible repository of data for subsequent analysis.
@@ 3.3 Creating the Open-set 3D Representation
Figure 3: Clustering in 3D of object instances using semantic embeddings and volumetric overlap. Semantic similarity is verified using CLIP and DINOv2 embeddings. Volumetric overlap is calculated using 3D bounding boxes and overlap matrices. The above figure shows the clustering process between two camera poses, represented by the coloured frustum, and the corresponding object point clouds are the same colour. For the O3D-SIM pipeline, this process is repeated for each camera pose, and the objects from each frame are either merged into an existing instance or added as a new instance into the incrementally built representation. The first example shows a positive case of merging, where a comparison is made for the same instance of a chair from different poses. The chairs are merged into a single instance due to success in both semantic similarity and volumetric overlap. The second example shows a case where the table (red) and the skateboard (blue), being very close to each other, have a volumetric overlap but are not merged due to a failure in semantic similarity, creating 2 different object instances. Example 3 shows an example of two separate chairs near each other. Since they are 2 different chairs, they fail to have a volumetric overlap, creating 2 separate instances.
@@ 3.3.1 The O3D-SIM Intialization
The set of object’s point clouds is processed further using DBSCAN[38] clustering for representation refinement. The point cloud is downsampled via voxel grid filtering to reduce the number of points and the computational complexity while preserving the data spatial structure manageable.
DBSCAN groups the points that are closely packed together while labelling points that lie alone in low-density regions as noise. In a post-clustering step, the largest cluster typically corresponds to the main object of interest within the point cloud is identified. This helps filter out the noise and irrelevant points, producing a cleaner representation of the object of interest.
@@ 3.3.2 Incremental Update of the O3D-SIM
Objects are detected for each new image, and new nodes are created like the initial image. These temporary nodes contain the 3D data for newly detected objects that must either be merged into the existing scene or added as new nodes. The similarity between newly detected and existing scene nodes is determined by combining visual similarity, derived from feature embeddings, and spatial similarity, obtained from point cloud overlap, to formulate an aggregate similarity measure. If this measure surpasses a predetermined threshold, the new detection is deemed to correspond to an existing object in the scene. Indeed, the newly detected node is either merged with an existing scene node or added as a new node.
The volumetric overlap of two point clouds is defined as the proportion of points in point cloud $pt_{i}$ that have nearest neighbours in point cloud $pt_{j}$, within a distance threshold $\delta_{nn}$. The metric can be mathematically expressed as:
$\text{Volumetric overlap}=\text{nnratio}(pt_{i},pt_{j})=\frac{1}{|pt_{i}|}\sum_{p\in pt_{i}}\mathbbm{1}(\min_{q\in pt_{j}}\|p-q\|\leq\delta_{nn})$ (1)
where $\mathbbm{1}$ is the indicator function that is 1 if the condition is true and 0 otherwise, $p$ and $q$ are points in point clouds $pt_{i}$ and $pt_{j}$, respectively, and $\|p-q\|$ is the Euclidean distance between points $p$ and $q$.
The semantic similarity between two point clouds is calculated using the cosine similarity between the DINO features of the image masks used to create the point clouds. This similarity is scaled to be between 0 and 1. It can be formulated as:
$\text{Semantic Similarity}=\frac{1}{2}\left(1+\frac{\mathbf{f}_{pt_{i}}\cdot\mathbf{f}_{pt_{j}}}{\|\mathbf{f}_{pt_{i}}\|\|\mathbf{f}_{pt_{j}}\|}\right)$ (2)
where $\mathbf{f}_{pt_{i}}$ and $\mathbf{f}_{pt_{j}}$ are the DINO feature vectors of the image masks corresponding to point clouds $pt_{i}$ and $pt_{j}$, respectively, $\cdot$ denotes the dot product, and $\|\cdot\|$ represents the norm of the vector.
@@ 3.4 Language-Guided Navigation
In this section, we leverage the LLM-based approach from [1], which uses ChatGPT 3.5 [39] to understand and map language commands to pre-defined function primitives that the robot can understand and execute. We have utilized the ChatGPT 3.5 API for our project due to its user-friendly interface and to have an accurate and fair comparison with SI Maps [1], as it also uses ChatGPT. It’s noteworthy that other large language models (LLMs), such as LLAMA [40], can achieve comparable results when provided with appropriate context for this process..
However, there are a few differences between our current approach and the approach in [1] regarding the use case of the LLM and the implementation of our function primitives. The previous approach used the LLM’s ability to bring in an open-set understanding by mapping general queries to the already-known closed-set class labels obtained via Mask2Former [7]. However, given the open-set nature of our new representation, O3D-SIM, the LLM does not need to do that. Figure  4 shows both approaches’ code output differences. The function primitives work similarly to the older approach, requiring the desired object type and its instance as an input. But now, the desired object is not from a pre-defined set of classes but a small query defining the object, so the implementation to find the desired location changes. We use the text and image-aligned nature of CLIP embeddings to find the desired object, where the input description is passed to the model, and its corresponding embedding is used to find the object in O3D-SIM. A cosine similarity is calculated between the embedding of the description and all the embeddings of our representation. These are ranked in an order based on the input query, and the desired instance is selected. This helps localise the required object instance from the language query.
@@
Once the object and the desired instance have been localised using the language command, a goal point is generated so the robot can perform autonomous navigation towards this object. We create a top-down occupancy grid map from the point cloud generated from RGB-D data and poses. Since the object is represented as a point cloud, we have multiple points for a single instance; hence, we calculate the object’s centroid. The obstacles in the occupancy map are inflated by the robot’s radius, and reachable free space is marked in this inflated map using Breadth-First Search (BFS). Finally, the object centroid is projected onto the grid map, and the closest reachable free grid cell to the centroid is returned to the robot as the goal point for autonomous navigation (Figure 5), therefore achieving Language-Guided Navigation.
Figure 5: This figure shows the process of generating a goal point once the target object has been localised based on a sample language query and O3D-SIM.
@@ 4 Experiments
Having introduced the O3D-SIM creation pipeline and its integration with ChatGPT for natural language understanding and Vision-Language Navigation (VLN) enhancement, we now turn to the evaluation of this novel representation both quantitatively and qualitatively. This will also shed light on the impact of the O3D-SIM representation on an agent’s ability to execute queries that mimic human interaction. Figure 1 shows a few examples of such queries. The evaluation is structured into two subsections: Section 4.2 focuses on the quantitative assessment of O3D-SIM, and Section 4.3 addresses the qualitative analysis of the representation. The section 4.1 details our hardware configurations for creating and testing the maps during real-world experiments.
4.1 Hardware Setup
Creation of O3D-SIM is an offline pipeline and leverages large foundational models that require a good amount of GPU VRAM. Hence, we use a server setup with a 16 GB Nvidia RTX A4000, 64 GB RAM and an AMD Ryzen 7 3800X 8-core processor to create the maps. For collecting data in the real world and testing out the navigation performance, we use a custom robot we have built in our lab, which provides wheel odometry using encoders, planar lidar scans using an RPLiDAR sensor and RGB-D data using a Kinect Azure Camera. For inference on the robot, we use a laptop with a 6 GB Nvidia RTX 3060, 32 GB RAM and an AMD Ryzen 7 5800H 8-core processor. We load O3D-SIM on the laptop and integrate it with the ROS Navigation Stack, where the goal is generated from language commands and passed to the mobile robot (section 3.4). We do not require the GPU on the laptop to infer from the CLIP model and O3D-SIM, and the object localisation using O3D-SIM and CLIP and the corresponding goal point generation happens in about a second. The inference time for the language query to code generation depends on the response time of the LLM’s API. It took an average time of about 45-60 seconds with ChatGPT 3.5 for our experiments, which is consistent with the time reported in other works.
@@ 4.2 Quantitative Evaluation
To facilitate the construction of O3D-SIM and its quantitative evaluation, we employ the Matterport3D dataset [41] within the Habitat simulator [42]. Matterport3D, a comprehensive RGB-D dataset, encompasses 10,800 panoramic views derived from 194,400 RGB-D images across 90 large-scale buildings. It offers surface reconstructions, camera poses, and 2D and 3D semantic segmentations — critical components for creating accurate Ground-truth models. Both Matterport3D and Habitat are widely utilized for assessing the navigational abilities of VLN agents in indoor settings, enabling robots to execute navigational tasks dictated by natural language commands in a seamless environment, with performance meticulously documented. To evaluate O3D-SIM, we compiled 5,267 RGB-D frames and their respective pose data from five distinct scenes, applying this dataset across all mapping pipelines included in our assessment. We also gathered real-world environment data for evaluation purposes, expanding our analysis to encompass six unique scenes.
@@
Results: We present the results of the evaluation metric Success Rate in Table 1. In our experimentation, we observe a remarkable improvement in performance compared to the other approaches we have shown in our paper, especially against the baselines from [1]. O3D-SIM performs better than VLMaps with CC due to its ability to identify instances robustly. SI Maps and O3D-SIM perform better than the baselines due to their ability to separate instances. However, O3D-SIM has the edge over SI Maps due to its open set and 3D nature, allowing it to understand the surroundings better. Figure 6 shows a few language queries we tested in our lab using all four representations - VLMaps, VLMaps_CC, SI Maps and O3D-SIM.
@@
Table 1: O3D-SIM outperforms the baseline methods from [1] by significantly large margins on the Success Rate metric. It also shows an improvement from our previous approach, i.e., SI-Maps. The best results are highlighted in bold. For this evaluation, the agent executes a set of open-set and closed-set queries in 5 different scenes from Matterport3D and 1 real world environment.
Figure 6: The figure shows the success/failure of the four representations in our lab environment for a few open/closed-set queries.
@@ 4.3 Qualitative Results
Our approach recognises instance-level semantics, accurately identifying 5 out of 6 table instances (with one false positive), underscoring its precision across both simulated and real-world data. This demonstrates the robustness of our pipeline, further evidenced by the clarity of the semantic map and the ease with which instance-level segmentation results can be visualized. While methods like VLMaps might identify a broader range of objects due to their open-set nature, and SI-Maps may detect multiple instances of the same object, O3D-SIM uniquely excels at both, offering a comprehensive solution.
Figure 7: This figure presents the qualitative results for O3D-SIM. Mapping sequence 1 was recorded using a Kinect Azure on a mobile robot, whereas mapping sequence 2 was recorded using an iPad, both in the real world. The environment was changed between the two sequences. The first four queries show some unique objects identified by O3D-SIM that normal closed-set approaches may fail to identify.